The calibration of a filling machine broke, causing it to underfill 5 jars of jam. The jars were accidentally mixed in with 13 properly filled jars. All the jars look the same. 3 jars are chosen from the batch and weighed. What is the probability that exactly 2 jars are underweight?
Round your answer to 3 decimal places.

Answer by stanbon(75887)

The calibration of a filling machine broke, causing it to underfill 5 jars of jam. The jars were accidentally mixed in with 13 properly filled jars. All the jars look the same. 3 jars are chosen from the batch and weighed. What is the probability that exactly 2 jars are underweight?
Round your answer to 3 decimal places.
---
Binomial Problem with n = 3 and p(under) = 5/18
---------------
P(x = 2) = 3C2(5/18)^2(13/18) = binompdf(3,5/18,2) = 0.1672
===============
Cheers,
Stan H.